Solve for x in this equation: x^2-5=8-2x^2
Answer by Boreal(15235)

x^2-5=8-2x^2
3x^2-13=0
3x^2=13
x^2=13/3
x=+/-sqrt(39)/3
Numerically 2.08
check
2.08^2-5=-0.67
8-2(2.08^2)=-0.65. These are equivalent given rounding. sqrt(39)/3 is the exact answer